OpenStudy (anonymous):

Val calculates the ratio of length and width of each section drawn on the grid and concludes that two sections are similar. Which two sections appear as similar geometric figures on the grid?

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (theeric):

Hi! I suggest you do what Val did! I suggest you look at each shape and divide the short side's length by the long side's length, and simplify! If you start with section H, you'd have the ratio as \(\dfrac{4}{6}=\dfrac{2}{3}\). Do that for every section until two ratios are equal. H: 2/3 K: S: T: M:

OpenStudy (anonymous):

I dont understand

OpenStudy (theeric):

Well, a ratio just compares two values by dividing one by the other. So I chose to do \(\dfrac{long\ side}{short\ side}\). Where did I lose you?

OpenStudy (anonymous):

Should I find the sides lengths with the distance formula firstt? I have 2 test to do todays. Chapter 5 and the segment one exam. So I'm really worried about getting a bad grade.

OpenStudy (theeric):

Good luck on both! To find the length of the sides, just count how many blocks it takes to get from one end to the other! :)

OpenStudy (anonymous):

Ah got it!

OpenStudy (anonymous):

so for K its 1/2 right?

OpenStudy (theeric):

My apologies, I said I would do short side divided by long, and changed that later. Keep it short divided by long because we already did that for H! K has a short side of two and a long side of three!

OpenStudy (theeric):

H: 2/3 K: S: T: M: So fill that in until two values are the same!

OpenStudy (theeric):

Have you found K? By the way, sorry for the delay! A pet of mine hit the escape key, which, as it turns out, deletes the entire response! So I had to retype it.

OpenStudy (anonymous):

I accidently just sent the exam in without answering this questions -.- I'm gonna move on to the next one. Thank you for helping me though

OpenStudy (theeric):

Sorry to hear that!! It's too bad that the program allows that... Not a good design... Good luck with the rest! It was H and K, by the way...

OpenStudy (anonymous):

I didn't even have a clue how to get it till you showed me the formula. So I apreciate it!